Please help me solve these two equations: 3x+4(1+2x)=15 and 4(1+3x)-2(3-x)=12.

Thank-you so much! I am revising for a test and preferably need it to be done step by step!

Answer by lwsshak3(11628)

3x+4(1+2x)=15 and 4(1+3x)-2(3-x)=12.
**
3x+4(1+2x)=15
clear parenthesis and distribute
3x+4+8x=15
add -4 to each side
3x+4-4+8x=15-4
combine like terms
11x=11
divide both sides by 11
x=1
..
4(1+3x)-2(3-x)=12
Clear parenthesis and distribute
4+12x-6+2x=12
combine like terms
14x-2=12
add 2 to both sides
14x=14
divide both sides by 14
x=1